OpenStudy (anonymous):

show that the graph of r = a cos θ + b sin θ is a circle, and find its center and radius

OpenStudy (anonymous):

multiply through by r to get \[r^2=ar\cos\theta +br\sin\theta\]

OpenStudy (anonymous):

covert to rectangular coordinates to get \[x^2+y^2=ax+by\]

OpenStudy (anonymous):

now get into the standard form for a circle by completing the square

OpenStudy (anonymous):

i.e. \[(x-\frac{a}{2})^2+(y-\frac{b}{2})^2=\frac{a^2+b^2}{4}\]

OpenStudy (anonymous):

so your center is \[(\frac{a}{2},\frac{b}{2})\] and the radius is \[\frac{a+b}{2}\]